OpenStudy (anonymous):

Rewrite x^2 - 6x + 9 as a square of a linear expression. Please Explain how to do this. a little lost

OpenStudy (anonymous):

It's just factoring, but evilly worded: (x-3)^2

OpenStudy (anonymous):

Oh, so is that the answer?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Please just a little bit of quick help please =)

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh yes, I see. That is the right answer, thanks so much for the clarification! =)

OpenStudy (anonymous):

It is (x-3)(x-3)

OpenStudy (solomonzelman):

Yes... And in any case, \(\large\color{slate}{\displaystyle x^2-ax+\left(\frac{a}{2} \right)^2 }\) you get, \(\large\color{slate}{\displaystyle (x-a)^2 }\)

OpenStudy (anonymous):

Thanks so much, for all your awesome help guys! I think i get it now =)